OpenStudy (anonymous):

Which of the following is a solution ? are solutions of the system 5x +2y-z=9 4x-2y+z=0 is it A) 1, 1 , -2 or B) 1 , -1, 2

OpenStudy (amistre64):

best way to test is, to try them out ....

OpenStudy (anonymous):

how can I test them out?

OpenStudy (amistre64):

you have place holders for values of x, y, and z. They give you values for x, y, and z. Replace the variables with the numbers, and see if you get a true solution for both equations.

OpenStudy (amistre64):

otherwise we would have to go thru a longer process of row reductions ...... and we would still be given parametric equations to test out

OpenStudy (nirmalnema):

put the values in place of x y and z..

OpenStudy (anonymous):

shouldn't there be 3 equations if there is 3 unknowns ?

OpenStudy (amistre64):

assuming this is linear algebra, the row reduction process would give us: http://www.wolframalpha.com/input/?i=rref%7B%7B5%2C2%2C-1%2C9%7D%2C%7B4%2C-2%2C1%2C0%7D%7D such that: x = 1 + 0 z y = 2 + 1/2 z z = 0 + 1 z

OpenStudy (nirmalnema):

yup @kelliegirl33 you are correct

OpenStudy (amistre64):

2 planes intersect in a line, at best

OpenStudy (anonymous):

could you just add them...that would eliminate the y's

OpenStudy (anonymous):

oh...and the z's

OpenStudy (amistre64):

x = 1 + 0 z x = 1 is good for both cases z = z is good for both cases y = 2 + 1/2 z if y = 1 z = 2(1-2) if y = -1 z = 2(-1-2)

OpenStudy (anonymous):

that explains alot :)

OpenStudy (amistre64):

:) but the row reduction by hand takes longer than the trial and error ... which is why i would have just gone with the plug and play method to start with

OpenStudy (anonymous):

that makes sense

OpenStudy (anonymous):

It is not posible ot be a.